The euler package

Frank Jensen^{∗}

1995/03/05

1 Introduction

The euler package provides a setup for using the AMS Euler family of fonts for math in L^{A}TEX documents. The AMS Euler family was designed by Hermann Zapf, commissioned by the American Mathematical Society. “The underlying philosophy of Zapf’s Euler design was to capture the flavor of mathematics as it might be written by a mathematician with excellent handwriting.” [2]

The euler package is based on Knuth’s macros for the book “Concrete Mathematics” [1]. Knuth’s macros can be found through anonymous ftp to labrea.

stanford.edu: look for the file gkpmac.tex in directory pub/tex/local/lib.

The Euler fonts can be found through anonymous ftp to e-math.ams.org: look

in directory pub/tex/amsfonts.

The purpose of the euler package is to provide the math part of the look of [1]. The other part (text fonts) is provided by the beton package. The reason for creating two packages is to make it easy to use the Euler math fonts together with other text fonts (in particular, it appears that the Euler fonts match many of the popular PostScript fonts pretty well).

Basically, the euler package provides the same setup (the same definitions, math codes, etc.) as gkpmac.tex with respect to the Euler fonts. However, Knuth [2] admits that the macros were written for one specific project, namely to typeset the “Concrete Mathematics” book [1]. So, the euler package actually does a little more than gkpmac.tex: for example, some ‘exotic’ symbols, present in CM math italic, are missing from the Euler fonts; the euler package takes care of this (whereas gkpmac.tex does not).

The use of the euler package requires a lot of care when entering the manuscript, since the package will cause, e.g., math numerals to come from the Euler Roman fonts; these numerals are easily distinguished from the normal text numerals. This implies that sloppy typing is clearly exposed: imagine how “$x$ is either 1 or $-1$” will appear when typeset! In general, the typist will in each case have to decide whether an input fragment is a math or a non-math entity. This holds true even for the tiniest details, such as punctuation characters (parentheses, colons, semicolons, commas, and periods).

Finally, a few words about the technical quality of the fonts (there are no doubts about their artistic quality): The fonts are not tuned for low resolution rendering: I recommend at least 600dpi. The fonts employ none of the tuning machinery of CM: e.g., they do not use the ‘blacker’ parameter specified in each Metafont mode definition, indicating that they might look somewhat light on printers that normally require a large ‘blacker’ value.

∗Updates by Frank Mittelbach.

2 What the euler package provides

The euler package defines two new math alphabet identifiers, \mathscr (Euler Script, uppercase letters only) and \mathfrak (Euler Fraktur, both lowercase and uppercase letters).

By default, the original meaning of \mathcal is preserved, but if the euler package is loaded with the mathcal option, \mathcal will produce Euler Script letters instead of the usual Computer Modern calligraphic letters.

Also, by default, the euler package does not redefine the \mathbf math alphabet identifier; it will still produce bold CM letters and digits. However, loading the euler package with the mathbf option will cause \mathbf to produce bold Euler Roman letters and digits. Another way to get bold symbols is to use the \boldsymbol command provided by the amsbsy package (part of AMS-L^{A}TEX).

Some users do not like the \hat accent glyph present in the Euler fonts (the euler package uses this accent glyph be default). To use the \hat accent glyph from the text font instead, load the euler package with the text-hat-accent option.

In [1], the space around relations in displays was increased (from ‘5mu plus 5mu’ to ‘10mu minus 3mu’). The euler package does not change this space, since this is (for most cases) controlled by the L^{A}TEX eqnarray environment. Also, in [1], displays are left justified with a suitable indentation; we leave such matters to the general layout.

References

[1] R. L. Graham, D. E. Knuth, and O. Patashnik. Concrete Mathematics. Addison-Wesley, 1989.

[2] D. E. Knuth. Typesetting Concrete Mathematics. TUGboat 10(1):31–36, 1989.